OpenStudy (anonymous):

solve for x. -3(x+1)-x/2+3/2<0

OpenStudy (amoodarya):

multiply by 2 -6(x+1)-x+3<0 -6x-6-x+3<0 -7x-3<0

OpenStudy (anonymous):

multiply -3 to (x+1) then it will become.. = -3x - 3 - x/2 + 3/2 < 0 =-3x - x/2 + 3/2 + 3 <0 =-7x/2 + 9/2 < 0 tranpose 9/2 then it will become negative.then solve for the value of x..